# Linearization in magnetoelasticity

### v1

@@ Abstract.
Starting from a model of nonlinear magnetoelasticity where magnetization is defined in the Eulerian configuration while elasticity is in the Lagrangean one, we rigorously derive a linearized model that coincides with the standard one that already appeared in the literature, see, e.g., [12] and where the zero-stress strain is quadratic in the magnetization. The relation of the nonlinear and linear model is stated in terms of the $\Gamma$-convergence and convergence of minimizers.
@@ 1. Introduction
Magnetoelasticity, also known as magnetostriction, is a characteristic observed in certain solids wherein there is a significant interdependence between mechanical and magnetic properties. This coupling manifests as reversible mechanical deformations that can be induced by applying an external magnetic field. This phenomenon holds considerable practical value, especially in the design of sensors, actuators, and various innovative functional-material devices, see e.g.  [17].
The origin of magnetoelasticity can be traced to the intricate interplay between the crystallographic structure of the material, where different crystals exhibit distinct easy axes of magnetization, and magnetic domains. In the absence of external magnetic fields, these magnetic domains align to minimize long-range dipolar effects, leading to a generally minimal or even negligible magnetization of the medium. Upon the application of an external magnetic field, the magnetic domains undergo a reorientation towards it. This reorientation leads to the emergence of a macroscopic deformation since magnetizations are linked to specific stress-free reference strains. As the magnetic field’s intensity increases, an increasing number of magnetic domains align themselves so that their principal axes of anisotropy align with the magnetic field in each region, eventually reaching saturation. For a deeper exploration of the foundational aspects of magnetoelasticity, refer to [7, 13].
Magnetoelasticity is a dynamic area of mathematical research reflecting modeling in linear and nonlinear regimes [12, 24] and analysis in the static and quasistatic setting [2, 5, 4] but also coupled with thermal and other phenomena, e.g.  [21]. Moreover, dimension-reduction problems covering lower-dimensional structures in various scaling regimes were recently analyzed, too; see, for instance, [4] or [11].
@@
(1.1) $\mathcal{G}(y,m):=\int_{\Omega}W(\nabla y,m\circ y){\,\rm d}x+\frac{1}{2}\int_{y(\Omega)}|\nabla m|^{2}{\,\rm d}y+\frac{\mu_{0}}{2}\int_{\mathbb{R}^{d}}|\nabla v_{m}|^{2}{\,\rm d}z$
for a deformation $y\in W^{1,p}(\Omega;\mathbb{R}^{d})$, $y=y_{0}$ on $\Gamma$, and a magnetization $m\in W^{1,2}(y(\Omega);\mathbb{R}^{d})$. In (1.1), the first term represents the elastic energy, the second term is the exchange energy, and the last term stands for magnetostatic energy; $\mu_{0}$ is the permittivity of void and $v_{m}$ is the magnetostatic potential generated by $m$. In particular, $v_{m}$ is a solution to the Maxwell equation $\nabla\cdot(-\mu_{0}\nabla v_{m}+\chi_{y(\Omega)}m)=0$ in $\mathbb{R}^{d}$ with $\chi_{y(\Omega)}$ the characteristic function of $y(\Omega)$. Furthermore, we assume that the specimen temperature is fixed and therefore the Heisenberg normalized constraint
@@
holds for all deformations.
@@
As in [14], we suppose the deformation gradient $F$ can be multiplicatively decomposed into magnetic and elastic parts $F=F_{m}F_{el}$, where the magnetic part $F_{m}$ is volume-preserving, i.e., $\det F_{m}=1$. In our setting, we assume $F_{m}:=\exp(-E(\mathcal{M}))$, i.e., $F^{-1}_{m}=\exp(E(\mathcal{M}))$, where $E(\mathcal{M})$ is defined as
$E(\mathcal{M}):=-\mathcal{M}\otimes\mathcal{M}+\frac{1}{d}I\,.$
@@
Here, $\mathcal{M}\in\mathbb{R}^{d}$ is a placeholder for the Lagrangean magnetization whose length $|\mathcal{M}|$ satisfies $\mathcal{M}\cdot\mathcal{M}=|\mathcal{M}|^{2}=1$, and $I$ is the identity matrix in $\mathbb{R}^{d\times d}$. Note that $\operatorname{tr}E(\mathcal{M})=0$, so that, indeed, $\det F_{m}^{-1}=1$. In view of the Heisenberg constraint (1.2), $M(x)=m(y(x))\det\nabla y(x)$ satisfies $|M(x)|=1$ for $x\in\Omega$. Consequently, if $\omega\subset\Omega$ we formally have the change-of-variables formula implying
$\int_{\omega}M(x)\,{\,\rm d}x=\int_{y(\omega)}m(z)\,{\,\rm d}z,$
@@
and $M$ and $m$ transform as volume densities in the reference (Lagrangean) and deformed (Eulerian) configurations, respectively.
Then we define
$e(F,m)=E(M)=-M\otimes M+\frac{1}{d}I=-(\det F)^{2}m\otimes m+\frac{1}{d}I\ .$
@@
In the case $F=I$, we simplify the notation by setting $e(I,m)=e(m):=-m\otimes m+\frac{1}{d}I$. Further, we assume that the elastic energy density is of form
$W(F,m)=\Phi\left(\exp(e(F,m))F\right)=\Phi(F_{el})\ ,$
@@
where $c>0$ is a material constant. We recall that the matrix exponential $\exp(A)=\sum_{k=0}^{\infty}A^{k}/k!$ is defined for every $A\in\mathbb{R}^{d\times d}$, setting also $A^{0}=I$. Here, $\Phi$ is a nonnegative function that is minimized at the identity.
@@
In the small-strain regime, we set $y_{\varepsilon}=id+\varepsilon u$ and assume that the influence of the magnetization on the elastic deformation of the specimen vanishes along $\varepsilon\to 0$ in the sense that $F_{m,\varepsilon}=\exp(\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\varepsilon E(M))$ and $\lim_{\varepsilon\to 0}F_{m,\varepsilon}=I$. Then, we consider the rescaled energy
(1.3) $\mathcal{G}_{\varepsilon}(u,m)$ $:=\frac{1}{\varepsilon^{2}}\int_{\Omega}W(Id+\varepsilon\nabla u(x),m(x+\varepsilon u(x))){\,\rm d}x+\frac{1}{2}\int_{y_{\varepsilon}(\Omega)}|\nabla m(y)|^{2}{\,\rm d}y+\frac{\mu_{0}}{2}\int_{\mathbb{R}^{d}}|\nabla v_{m}(z)|^{2}{\,\rm d}z$
$=\frac{1}{\varepsilon^{2}}\Phi(\exp(\varepsilon E(M))(I+\varepsilon\nabla u(x)))\,{\,\rm d}x\frac{1}{2}\int_{y_{\varepsilon}(\Omega)}|\nabla m(y)|^{2}{\,\rm d}y+\frac{\mu_{0}}{2}\int_{\mathbb{R}^{d}}|\nabla v_{m}(z)|^{2}{\,\rm d}z.$
@@
where $e(u):=\frac{1}{2}(\nabla u^{T}+\nabla u)$ is the symmetric part of the displacement gradient and $\mathbb{C}:=D^{2}\Phi(I)$ is the elasticity tensor. Note that (1.4) suggests a decomposition of the total strain $\bm{\epsilon}(u)$ into the magnetic part $-e(m)=m\otimes m-I/d$ and the elastic part $\bm{\epsilon}(u)+e(m)$. This decomposition was already used, e.g., in [12, 19, 20]. The aim of this contribution is to derive (1.4) rigorously by means of $\Gamma$-convergence.
@@
    $\Phi(F)\geq C\bigg{(}\frac{1}{(\det F)^{a}}-1\bigg{)}\qquad\text{for every $F\in\mathbb{M}^{d}$.}$
Frame-indifference (a) implies that $W(F,m)=W(F,-m)=W(QF,Qm)$ for all $F\in\mathbb{R}^{d\times d}$ all $Q\in\textrm{SO}(d)$, and all $m\in\mathbb{R}^{d}$. Moreover, for every $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}F\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\in\mathbb{R}^{d\times d}$
(1.6) $D^{2}\Phi(I)[\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}F,F\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}]=D^{2}\Phi(I)[\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}F_{sym},F_{sym}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}].$
Together with (d), it implies that $D^{2}\Phi(I)[\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}F,F\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}]=0$ if $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}F=-F^{\top}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}$ and
(1.7) $D^{2}\Phi(I)[\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}F_{sym},F_{sym}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}]\geq|\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}F_{sym}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}|^{2}.$
@@
In what follows, the work of a mechanical load is modeled by a continuous linear functional $\mathscr{L}\colon L^{2}(\Omega;\mathbb{R}^{d})\to\mathbb{R}$ and the magnetic loading is described by a functional $\mathscr{M}\colon L^{2}(\Omega;\mathbb{R}^{d})\times L^{2}(\mathbb{R}^{d};\mathbb{R}^{d})\to\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}$
@@
under the prescribed boundary conditions. We are interested in the case where the load has the form $\varepsilon\mathscr{L}$, and we want to study the behavior of the solution as $\varepsilon$ tends to zero. We write
$y=x+\varepsilon u$
and we assume the Dirichlet boundary condition of the form
$v=x+\varepsilon w\quad\mathcal{H}^{n-1}\text{-a.e. on }\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Gamma,\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}$
@@
$W^{1,2}_{w}(\Omega;\mathbb{R}^{d}):=\left\{u\in W^{1,2}(\Omega;\mathbb{R}^{d}):u=w\quad\mathcal{H}^{n-1}\text{-a.e. on }\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Gamma\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\right\}\!,$
the corresponding minimum problem for $u$ becomes
@@
$(u,m)\in W^{1,p}_{w}(\Omega;\mathbb{R}^{d})\times W^{1,2}(y_{\varepsilon}(\Omega);\mathbb{R}^{d})\biggr{\}}\!.$
@@
$\chi_{y_{\varepsilon}(\Omega)}m\in W^{1,2}(y_{\varepsilon}(\Omega),\mathbb{R}^{d}),\>|m\circ y_{\varepsilon}|\,\det\nabla y_{\varepsilon}=1\text{ a.e.~{}in }\Omega\bigr{\}},$
@@
Assume that $W\colon\mathbb{M}^{d}\times\mathbb{R}^{d}\rightarrow[0,\infty]$ satisfies conditions (a)(e) for some $p>d$, and let $w\in W^{2,\infty}(\Omega;\mathbb{R}^{d})$. For every $\varepsilon_{j}\to 0$ we have that
@@
Under the hypotheses of Theorem 1.1, for every $\varepsilon>0$ define $s_{\varepsilon}:=\inf\{\mathcal{F}_{\varepsilon}(u,m):(u,m)\in\mathcal{A}^{w}_{\varepsilon}\}$ and let $\left\{(u_{\varepsilon},m_{\varepsilon})\right\}_{\varepsilon>0}$ be a sequence such that
$\mathcal{F}_{\varepsilon}(u_{\varepsilon},m_{\varepsilon})=s_{\varepsilon}+o(1).$
Then, $u_{\varepsilon}\rightharpoonup u_{0}$ weakly in $W^{1,2}(\Omega;\mathbb{R}^{d})$, $\chi_{y_{\varepsilon}(\Omega)}m_{\varepsilon}\to\chi_{\Omega}m_{0}$ in $L^{2}(\mathbb{R}^{d};\mathbb{R}^{d})$, and $\chi_{y_{\varepsilon}(\Omega)}\nabla m_{\varepsilon}\rightharpoonup\chi_{\Omega}\nabla m_{0}$ weakly in $L^{2}(\mathbb{R}^{d};\mathbb{R}^{d})$, where $(u_{0},m_{0})\in\mathcal{A}_{0}^{w}$ is the solution to the minimum problem
@@
Under the hypotheses of Theorem 1.1, $(u_{\varepsilon},m_{\varepsilon})\to(u_{0},m_{0})$ strongly in $W^{1,2}(\Omega;\mathbb{R}^{d})\times L^{2}(\mathbb{R}^{d};\mathbb{R}^{d})$.
The proof of Theorems 1.11.3 is postponed to Sections 4. Compactness issues are discussed in Section 3.

### v2

@@ Abstract.
Starting from a model of nonlinear magnetoelasticity where magnetization is defined in the Eulerian configuration while elastic deformation is in the Lagrangian one, we rigorously derive a linearized model that coincides with the standard one that already appeared in the literature, see, e.g., [19] and where the zero-stress strain is quadratic in the magnetization. The relation of the nonlinear and linear model is stated in terms of the $\Gamma$-convergence and convergence of minimizers.
@@ 1. Introduction
Magnetoelasticity is a characteristic observed in certain solids wherein there is a significant interdependence between their mechanical and magnetic properties. This coupling manifests as reversible mechanical deformations that can be induced by applying an external magnetic field. This phenomenon holds considerable practical value, especially in the design of sensors, actuators, and various innovative functional-material devices, see e.g.  [28].
The origin of magnetoelasticity can be traced to the intricate interplay between the crystallographic structure of the material, where different crystals exhibit distinct easy axes of magnetization, and magnetic domains [31]. In the absence of external magnetic fields, these magnetic domains align to minimize long-range dipolar effects, leading to a generally minimal or even negligible magnetization of the medium. Upon the application of an external magnetic field, the magnetic domains undergo a reorientation towards it. This reorientation leads to the emergence of a macroscopic deformation since magnetizations are linked to specific stress-free reference strains. As the magnetic field’s intensity increases, an increasing number of magnetic domains align themselves so that their principal axes of anisotropy align with the magnetic field in each region, eventually reaching saturation. For a deeper exploration of the foundational aspects of magnetoelasticity, refer to [14, 23, 31].
Magnetoelasticity is a dynamic area of mathematical research which reflects modeling in linear and nonlinear regimes [19, 40], analysis in the static and quasistatic setting [5, 12, 10] while also being coupled with thermal and other phenomena, e.g.  [37]. Moreover, dimension-reduction problems covering lower-dimensional structures in various scaling regimes were recently as well analyzed; for instance, see [10] or [18].
@@
(1.1) $\mathcal{G}(y,m):=\int_{\Omega}W(\nabla y,m\circ y){\,\rm d}x+\frac{1}{2}\int_{y(\Omega)}|\nabla m|^{2}{\,\rm d}z+\frac{\mu_{0}}{2}\int_{\mathbb{R}^{d}}|\nabla v_{m}|^{2}{\,\rm d}z$
for a deformation $y\in W^{1,p}(\Omega;\mathbb{R}^{d})$ with $p>d$, $\det\nabla y>0$ a.e., $y$ is injective a.e. in $\Omega$, and $y=y_{0}$ on $\Gamma$ where $y_{0}\in W^{1,p}(\Omega;\mathbb{R}^{d})$ is a given boundary datum, and a magnetization $m\in W^{1,2}(y(\Omega)\setminus y(\partial\Omega);\mathbb{R}^{d})$. In (1.1), the first term represents the elastic energy, the second term is the exchange energy, and the last term stands for magnetostatic energy; $\mu_{0}$ is the permittivity of void and $v_{m}$ is the magnetostatic potential generated by $m$. In particular, $v_{m}$ is a solution to the Maxwell equation $\nabla\cdot(-\mu_{0}\nabla v_{m}+\chi_{y(\Omega)}m)=0$ in $\mathbb{R}^{d}$ with $\chi_{y(\Omega)}$ the characteristic function of $y(\Omega)$. Furthermore, we can assume that the specimen temperature is fixed and therefore the Heisenberg normalized constraint
@@
holds for all deformations [14, 32].
Let us point out some peculiarities caused by the mixed Eulerian–Lagrangian formulation. First, to correctly define the composition $m\circ y$, we must identify $y$ with its continuous representative, due to Sobolev embeddings theorems [34, Section 1.4.5]. Second, as it is conventional in magnetoelasticity [39, 12, 13], we define $m\in W^{1,2}(y(\Omega)\setminus y(\partial\Omega);\mathbb{R}^{d})$, since $y(\Omega)\setminus y(\partial\Omega)$ is an open set and it differs form $y(\Omega)$ just on a set of measure zero, i.e. $|y(\Omega)|=|y(\Omega)\setminus y(\partial\Omega)|$, see for example [25] or [12, Lemma 2.1].
@@
As in [24, 41], we suppose the deformation gradient $F$ can be multiplicatively decomposed into magnetic and elastic parts $F=F_{m}F_{el}$ with the magnetic part $F_{m}$ being volume-preserving, i.e., $\det F_{m}=1$. In our setting, we assume $F_{m}:=\exp(-E(M))$, i.e., $F^{-1}_{m}=\exp(E(M))$, where $E(M)$ is defined as
$E(M):=-M\otimes M+\frac{1}{d}I.$
@@
Here, $M\in\mathbb{R}^{d}$ is a placeholder for the Lagrangian magnetization whose length $|M|$ satisfies $M\cdot M=|M|^{2}=1$, and $I$ is the identity matrix in $\mathbb{R}^{d\times d}$. Note that $\operatorname{tr}E(M)=0$, so that, indeed, $\det F_{m}^{-1}=1$. In view of the Heisenberg constraint (1.2),
(1.3) $M(x)=m(y(x))\det\nabla y(x)$
satisfies $|M(x)|=1$ for $x\in\Omega$. Consequently, using the change-of-variables formula for Sobolev mappings [29] (considering that $y$ is a.e. injective and satisfies the Lusin $(N)$-condition) for $\omega\subset\Omega$ it holds
$\int_{\omega}M(x)\,{\,\rm d}x=\int_{y(\omega)}m(z)\,{\,\rm d}z.$
@@
Therefore, $M$ and $m$ transform as volume densities in the reference (Lagrangian) and the deformed (Eulerian) configurations, respectively.
We define $e\colon\mathbb{R}^{d\times d}\times\mathbb{R}^{d}\to\mathbb{R}^{d\times d}$ as
$e(F,m)=-(\det F)^{2}m\otimes m+\frac{1}{d}I.$
Whenever $m$ and $M$ are two magnetic fields related by (1.3), then $e(\nabla y,m\circ y)=E(M)$ in $\Omega$.
@@
In the case $F=I$, we simplify the notation by setting $e(m):=e(I,m)=-m\otimes m+\frac{1}{d}I$. Next, we assume that the elastic energy density is of form
$W(F,m)=\Phi\left(\exp(e(F,m))F\right)=\Phi(F_{el}).$
@@
We recall that the matrix exponential $\exp(A)=\sum_{k=0}^{\infty}A^{k}/k!$ is defined for every $A\in\mathbb{R}^{d\times d}$, setting also $A^{0}=I$. Here, $\Phi$ is a nonnegative function that is minimized precisely at the set of proper rotations. This form of the energy density is inspired by energy expressions in the following papers on nematic elastomers [1, 2, 3, 20].
@@
In the small-strain regime, we set $y_{\varepsilon}=id+\varepsilon u$ and assume that the influence of the magnetization on the elastic deformation of the specimen vanishes along $\varepsilon\to 0$ in the sense that $F_{m,\varepsilon}=\exp(-\varepsilon E(M))$ and $\lim_{\varepsilon\to 0}F_{m,\varepsilon}=I$. Then, we consider the rescaled energy
(1.4) $\mathcal{G}_{\varepsilon}(u,m)$ $:=\frac{1}{\varepsilon^{2}}\int_{\Omega}W(Id+\varepsilon\nabla u(x),m(x+\varepsilon u(x))){\,\rm d}x+\frac{1}{2}\int_{y_{\varepsilon}(\Omega)}|\nabla m(z)|^{2}{\,\rm d}z+\frac{\mu_{0}}{2}\int_{\mathbb{R}^{d}}|\nabla v_{m}(z)|^{2}{\,\rm d}z$
$=\frac{1}{\varepsilon^{2}}\int_{\Omega}\Phi(\exp(\varepsilon E(M))(I+\varepsilon\nabla u(x)))\,{\,\rm d}x+\frac{1}{2}\int_{y_{\varepsilon}(\Omega)}|\nabla m(z)|^{2}{\,\rm d}z+\frac{\mu_{0}}{2}\int_{\mathbb{R}^{d}}|\nabla v_{m}(z)|^{2}{\,\rm d}z.$
@@
where $e(u):=\frac{1}{2}(\nabla u^{T}+\nabla u)$ is the symmetric part of the displacement gradient and $\mathbb{C}:=D^{2}\Phi(I)$ is the elasticity tensor. Note that (1.5) suggests a decomposition of the total strain $\bm{\epsilon}(u)$ into the magnetic part $-e(m)=m\otimes m-I/d$ and the elastic part $\bm{\epsilon}(u)+e(m)$. This decomposition has already been used, e.g., in [19, 33, 35]. The aim of this contribution is to derive (1.5) rigorously by means of $\Gamma$-convergence.
@@
    $\Phi(F)\geq C\bigg{(}\frac{1}{(\det F)^{a}}-1\bigg{)}\qquad\text{for every $F\in{\mathbb{R}^{d\times d}}$.}$
Let us notice that $g_{p}$ satisfies the following properties: $g_{p}$ is convex,
(1.7) $g_{p}(t)\leq\frac{1}{p}\min\{t^{p},t^{2}\}\,,\qquad g_{p}(s+t)\leq C(g_{p}(s)+t^{2})$
for some constant $C>0$ depending on $p$, and for any $C_{1}>0$ there exists a constant $C_{2}>0$, depending only on $p$ and on $C_{1}$, such that
(1.8) $g_{p}(C_{1}t)\leq C_{2}g_{p}(t)\,.$
Frame-indifference (a) implies that $W(F,m)=W(F,-m)=W(QF,Qm)$ for all $F\in\mathbb{R}^{d\times d}$ all $Q\in\textrm{SO}(d)$, and all $m\in\mathbb{R}^{d}$. Moreover, for every $F\in\mathbb{R}^{d\times d}$
(1.9) $D^{2}\Phi(I)[F,F]=D^{2}\Phi(I)[F_{sym},F_{sym}].$
Together with (d), it implies that $D^{2}\Phi(I)[F,F]=0$ if $F=-F^{\top}$ and
(1.10) $D^{2}\Phi(I)[F_{sym},F_{sym}]\geq|F_{sym}|^{2}.$
@@
In what follows, the work of a mechanical load is modeled by a continuous linear functional $\mathscr{L}\colon L^{2}(\Omega;\mathbb{R}^{d})\to\mathbb{R}$ and the magnetic loading is described by a functional $\mathscr{M}\colon L^{2}(\Omega;\mathbb{R}^{d})\times L^{2}(\mathbb{R}^{d};\mathbb{R}^{d})\to\mathbb{R}$
@@
under the prescribed boundary conditions. The area of interest of this paper is the case where the load has the form $\varepsilon\mathscr{L}$, and to study the behavior of the solution as $\varepsilon$ tends to zero. We write
$y_{\varepsilon}=id+\varepsilon u$
and assume the Dirichlet boundary condition of the form
$y_{\varepsilon}=id+\varepsilon w\quad\mathcal{H}^{d-1}\text{-a.e. on }\Gamma,$
@@
$W^{1,p}_{w}(\Omega;\mathbb{R}^{d})$ ::=:= $\left\{u\in W^{1,p}(\Omega;\mathbb{R}^{d}):u=w\quad\mathcal{H}^{d-1}\text{-a.e. on }\Gamma\right\}\!,$
$W^{2,\infty}_{w}(\Omega;\mathbb{R}^{d})$ ::=:= $\left\{u\in W^{2,\infty}(\Omega;\mathbb{R}^{d}):u=w\quad\mathcal{H}^{d-1}\text{-a.e. on }\Gamma\right\}\!,$
the corresponding minimum problem for $(u,m)$ becomes
@@
$(u,m)\in W^{1,p}_{w}(\Omega;\mathbb{R}^{d})\times W^{1,2}(y_{\varepsilon}(\Omega)\setminus y_{\varepsilon}(\partial\Omega);\mathbb{R}^{d})\biggr{\}}\!.$
@@
$m\in W^{1,2}(y_{\varepsilon}(\Omega)\setminus y_{\varepsilon}(\partial\Omega);\mathbb{R}^{d}),\>|m\circ y_{\varepsilon}|\,\det\nabla y_{\varepsilon}=1\text{ a.e.~{}in }\Omega\bigr{\}},$
@@
Assume that $W\colon\mathbb{R}^{d\times d}\times\mathbb{R}^{d}\rightarrow[0,\infty]$ satisfies conditions (a)(e) for some $p>d$, and let $w\in W^{2,\infty}(\Omega;\mathbb{R}^{d})$. For every $\varepsilon_{j}\to 0$ we have that
@@
Remark 1.2.
In view of the compactness properties stated in Proposition 3.1 and in order to state the $\Gamma$-convergence in $W^{1,2}(\Omega;\mathbb{R}^{d})\times L^{2}(\mathbb{R}^{d};\mathbb{R}^{d})$, In Theorem 1.1 we identify $m_{\varepsilon}$ with $\chi_{y_{\varepsilon}(\Omega)}m_{\varepsilon}$ for $(u_{\varepsilon},m_{\varepsilon})\in\mathcal{A}^{w}_{\varepsilon}$ and $y_{\varepsilon}=id+\varepsilon u_{\varepsilon}$. Similarly, in the limit as $\varepsilon\to 0$ we identify the magnetization $m$ with $\chi_{\Omega}m$.
@@
Under the hypotheses of Theorem 1.1, for every $\varepsilon>0$ define $s_{\varepsilon}:=\inf\{\mathcal{F}_{\varepsilon}(u,m):(u,m)\in\mathcal{A}^{w}_{\varepsilon}\}$ and let $(u_{\varepsilon},m_{\varepsilon})\in\mathcal{A}^{w}_{\varepsilon}$ be a sequence such that
(1.13) $\mathcal{F}_{\varepsilon}(u_{\varepsilon},m_{\varepsilon})=s_{\varepsilon}+o(1)\quad\text{ as }\varepsilon\to 0.$
Then, $u_{\varepsilon}\rightharpoonup u_{0}$ weakly in $W^{1,2}(\Omega;\mathbb{R}^{d})$, $\chi_{y_{\varepsilon}(\Omega)}m_{\varepsilon}\to\chi_{\Omega}m_{0}$ in $L^{2}(\mathbb{R}^{d};\mathbb{R}^{d})$, and $\chi_{y_{\varepsilon}(\Omega)}\nabla m_{\varepsilon}\rightharpoonup\chi_{\Omega}\nabla m_{0}$ weakly in $L^{2}(\mathbb{R}^{d};\mathbb{R}^{d\times d})$, where $(u_{0},m_{0})\in\mathcal{A}_{0}^{w}$ is the solution to the minimum problem
@@
Under the hypotheses of Theorems 1.1 and 1.3, let $(u_{\varepsilon},m_{\varepsilon})\in\mathcal{A}^{w}_{\varepsilon}$ satisfy (1.13). Then, up to a subsequence, $(u_{\varepsilon},\chi_{y_{\varepsilon}(\Omega)}m_{\varepsilon})\to(u_{0},\chi_{\Omega}m_{0})$ strongly in $W^{1,2}(\Omega;\mathbb{R}^{d})\times L^{2}(\mathbb{R}^{d};\mathbb{R}^{d})$.
The proof of Theorems 1.11.4 is postponed to Section 4. Compactness issues are discussed in Section 3.
Remark 1.5 (Discussion for the case $p\leq d$).
Let us notice that the $p$-integrability of deformations’ gradients plays an essential role only for the compactness. The result of this paper can be extended for the case $p=d$ using the means of quasiconformal analysis, see details in, e.g., [30]. First, an admissible deformation $y\in\mathcal{Y}$ is continuous and differentiable a.e., and a continuous representative satisfies the Lusin $(N)$-condition. Thus, the composition $m\circ y$ is well defined, and the change-of-variables formula is valid for the chosen representative. Second, instead of $y(\Omega)\setminus y(\partial\Omega)$ one needs to use the topological image, see e.g. [36, 5], to define $m$ properly. Finally, there is no uniform convergence $y_{\varepsilon}\to y$, only locally uniform convergence. Therefore, the proof of compactness in Proposition 3.1 needs more refined estimates.
There are several options in the literature, which offer ways to deal with the case $p>d-1$. For instance, one may use the approach prohibiting cavitations as in Müller–Spector [36], Barchiesi–Henao–Mora-Coral [5], and Bresciani [11]. Although a stronger coercivity condition is needed to guarantee weak convergence of determinants. Another option, is to consider deformations with finite surface energy as in [5], (INV)-condition or weak limits of homeomorphisms [9, 21]. Admissible deformations in the case $p=d-1$ may be considered within the limits of homeomorphisms [22] or mappings with finite surface energy [6, 7].